12 quarts of an acid solution was mixed with 3 quarts of pure water to make it 52% solution. Find the percent concentration of the first solution.
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

12 quart of x solution mixed with 3 quarts of 1 solution (1 is 100%, it's 100% water)
So our total will be 12+3 = 15 and we want a 0.52 solution
:
12x+3(1) = 15(0.52)
12x+3 = 7.8
12x = 4.8
x = 0.4
the 12 quarts are a 40% solution

Answer by ikleyn(52787)

.
12 quart of x solution mixed with 3 quarts of pure water (0 is 0%, it's 0% acid solution).
So our total will be 12+3 = 15 and we want a 0.52 solution
:
12x + 3*0 = (12+3)*0.52
12x = 7.8
x = 0.65
the 12 quarts are a 65% solution.
------------------

Solved. The solution by the other tutor is wrong. It produces an "impossible" answer.


There is entire bunch of lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for alloys
    - Typical word problems on mixtures from the archive
    - OVERVIEW of lessons on word problems for mixtures
in this site.

Read them and become an expert in solution the mixture word problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Mixture problems